# Towards a test of the Born rule in high-energy collisions

### v2

@@ 4 Testing tau leptons with average helicity measurements
Figure 7: Feynman diagram of $Z$-boson production and decay to $\tau^{+}\tau^{-}$ in a $pp$ collision. Quarks from the colliding protons ($q$ and $\overline{q}$) annihilate to produce a $Z$-boson, which subsequently decays into a $\tau^{+}\tau^{-}$ pair. Each $\tau$ then decays into a tau neutrino ($\nu_{\tau}$ or $\overline{\nu}_{\tau}$) and a charged pion ($\pi^{-}$ or $\pi^{+}$).
To produce $\tau$ leptons, $pp$ collisions at the LHC are typically used (see for example Fig. 7 for a quark-antiquark annihilation into $Z$). Both CMS [63] and ATLAS [64] collect millions of $Z\rightarrow\tau\tau$ events with integrated luminosities above $100\,\mathrm{fb}^{-1}$. In this study, we focus on $\tau^{+}\tau^{-}$ pairs originating from $Z$ bosons. Because $pp$ collisions produce a wide range of final states, one must distinguish genuine $\tau$ leptons from other processes such as top-quark or $W$-boson decays. The procedure is analogous to how $\pi^{0}$ mesons are identified by their two-photon mass near $135\,\mathrm{MeV}$. Here, rather than looking for a $\gamma\gamma$ peak, we use specialised $\tau$-identification algorithms to reconstruct the $\tau^{+}\tau^{-}$ invariant mass near the $Z$-boson peak ($\sim 91\,\mathrm{GeV}$). Events lying outside a suitable mass window (such as ±$\pm 3\sigma$) are classified as background.
We may mention that a recent study has looked at reconstructing the complete density operator $\hat{\rho}$ for $\tau^{+}\tau^{-}$ pairs produced in $e^{+}e^{-}$ annihilation, and has used the resulting $\tau$‑polarisation observables to probe physics beyond the Standard Model (for example, setting limits on the $\tau$ electric dipole moment) [65]. However, because this production channel does not originate from an especially short-timescale process, here we have instead focused on $\tau$ leptons originating from $Z$ decays. Also, instead of reconstructing the complete mean polarisation vector $\mathbf{P}$ (which for our two-state system is equivalent to $\hat{\rho}$), we have limited ourselves to reconstructing what is in effect the longitudinal component of $\mathbf{P}$ (the mean helicity). More recently, polarisations reconstructed from scattering kinematics have also been employed to test another fundamental feature of quantum theory, namely contextuality [66, 67]. Some of that work also looks at polarisation for $W$-bosons created by (very short timescale) top-quark decays, and therefore may have some bearing on the physics explored here.
@@ 4.1 Pure-state formula at fixed kinematics
In a highly idealised situation, namely a single kinematic configuration at fixed centre-of-mass energy $\sqrt{s}$ and fixed scattering angle $\theta$, the spin state of the $\tau^{-}\tau^{+}$ pairs can be written in amplitude form
$|\tau\rangle=\;\mathcal{M}_{+}\,\lvert+-\rangle\;+\;\mathcal{M}_{-}\,\lvert-+\rangle,$ (29)
@@
Let it be clear that |$\lvert$+$\rangle$| + - ⟩ refers to $\tau^{-}$ with helicity $+\tfrac{1}{2}$ and $\tau^{+}$ with helicity $-\tfrac{1}{2}$, whereas |$\lvert$-$\rangle$| - + ⟩ is the opposite. Thus, $\mathcal{M}_{+}$ corresponds to $\tau^{-}$ having helicity $+\tfrac{1}{2}$, and $\mathcal{M}_{-}$ corresponds to $\tau^{-}$ having helicity $-\tfrac{1}{2}$. No summation over angles or energies is included. In a perfect one-angle, one-energy scenario, such as an $e^{+}e^{-}$ collision at a fixed $\sqrt{s}$ and fixed $\theta$, this amplitude-level expression is essentially the mean helicity of the sub-process.
@@ 4.2 Ensemble-level cross-section formula
[…]
4.3 Extracting mean tau helicity experimentally
@@
where $I$ is the incident flux and $N$ is the observed number of events within a given kinematic configuration. Counting events in different angular bins thus serves as a proxy for the underlying cross section.
In the absence of detector effects and assuming full knowledge of the $\tau$ rest frame, the two-body decay kinematics imply that the direction of the visible decay product is determined by the spin of the parent $\tau$, due to angular momentum conservation and the fixed spin states of the daughter particles. For $\tau^{-}\rightarrow\pi^{-}\nu_{\tau}$, the neutrino is always left-handed (with spin opposite to its momentum). Therefore
  • A $\tau^{-}$ with spin $+\tfrac{1}{2}$ emits the $\pi^{-}$ along the spin direction ($\cos\theta>0$);
  • A $\tau^{-}$ with spin $-\tfrac{1}{2}$ emits the $\pi^{-}$ opposite to the spin direction ($\cos\theta<0$).
Here, $\theta$ is the angle between the $\tau$ spin direction and the pion momentum. Even though we are considering the $\tau$ is in its rest frame and helicity is not strictly defined, the pion direction remains directly correlated with the spin state of the $\tau$. In this way, the pion emission angle serves as an experimental proxy for helicity.
@@
where $N_{f}$ is the number of events in the forward region and $N_{b}$ is the number in the backward region. Because the helicity state is (ideally) in one-to-one correspondence with the emission direction of the pion, this ratio captures the asymmetry of the $\tau^{-}$ with respect to $+\tfrac{1}{2}$ and $-\tfrac{1}{2}$ helicity.
@@ 5 Background assumptions and Born-rule deviations
In principle, there is clearly a tension between searching for deviations from the Born rule with an experiment whose analysis, in part, depends on the Born rule. However, in this paper we are concerned with testing the Born rule only as applied to simple two-state systems (spin or polarisation) whereas the experimental modelling and event selection applies the Born rule to multi-particle scattering amplitudes. On this point, there are two points to make. First, it could happen that the Born rule is violated only in the first case and not in the second case. Second, because multi-particle scattering observables are averaged over broad regions of phase space, it seems plausible that small deviations from the Born rule could be washed out, leaving significant deviations only for the simple two-state systems. While this issue deserves further consideration, on this basis we propose the following working hypothesis: we assume that the Born rule holds for background modelling and event selection, while allowing for possible deviations only in the two-state polarisation observables we propose to measure.
@@ 6 Conclusion
We have discussed how certain experiments at high-energy colliders might be repurposed as tests of the Born rule on timescales as short as $\sim 10^{-25}\,\mathrm{s}$. In ideal conditions, measuring the polarisation of individual gamma-ray photons would enable us to probe polarisation probabilities as a function of angle, providing detailed constraints on Born-rule violations, as parameterised here in terms of expectation values $E(\mathbf{m})$ that are nonlinear in the Bloch measurement axis $\mathbf{m}$. Unfortunately, for such high-energy photons, current technology appears to permit measurement of average polarisation only, limiting the search for Born-rule violations to anomalies in the effective polarisation vector $\mathbf{P}$, which need not coincide with the usual quantum vector $\mathbf{P}_{\mathrm{QT}}$ as defined by (5). It may be hoped that technical developments will eventually permit polarisation measurements for single gamma-ray photons, though this is a matter for future work.
@@
Acknowledgement. We warmly thank Oliver K. Baker for his interest in and support of this work.

### v4

@@ 4 Testing tau leptons with average helicity measurements
Figure 7: Feynman diagram of $Z$-boson production and decay to $\tau^{+}\tau^{-}$ in a pppp collision. Quarks from the colliding protons ($q$ and $\overline{q}$) annihilate to produce a $Z$-boson, which subsequently decays into a $\tau^{+}\tau^{-}$ pair. Each $\tau$ then decays into a tau neutrino ($\nu_{\tau}$ or $\overline{\nu}_{\tau}$) and a charged pion ($\pi^{-}$ or $\pi^{+}$).
To produce $\tau$ leptons, $pp$ collisions at the LHC are typically used (see for example Fig. 7 for a quark-antiquark annihilation into $Z$). Both CMS [65] and ATLAS [66] collect millions of $Z\rightarrow\tau\tau$ events with integrated luminosities above $100\,\mathrm{fb}^{-1}$. In this study, we focus on $\tau^{+}\tau^{-}$ pairs originating from $Z$ bosons. Because $pp$ collisions produce a wide range of final states, one must distinguish genuine $\tau$ leptons from other processes such as top-quark or $W$-boson decays. The procedure is analogous to how $\pi^{0}$ mesons are identified by their two-photon mass near $135\,\mathrm{MeV}$. Here, rather than looking for a $\gamma\gamma$ peak, we use specialised $\tau$-identification algorithms to reconstruct the $\tau^{+}\tau^{-}$ invariant mass near the $Z$-boson peak ($\sim 91\,\mathrm{GeV}$). Events lying outside a suitable mass window (such as $\pm 3\sigma$) are classified as background.
We may mention that a recent study has looked at reconstructing the complete density operator $\hat{\rho}$ for $\tau^{+}\tau^{-}$ pairs produced in $e^{+}e^{-}$ annihilation, and has used the resulting $\tau$‑polarisation observables to probe physics beyond the Standard Model (for example, setting limits on the $\tau$ electric dipole moment) [67]. However, because this production channel does not originate from an especially short-timescale process, here we have instead focused on $\tau$ leptons originating from $Z$ decays. Also, instead of reconstructing the complete mean polarisation vector $\mathbf{P}$ (which for a quantum two-state system is equivalent to $\hat{\rho}$), we have limited ourselves to reconstructing what is in effect the longitudinal component of $\mathbf{P}$ (the mean helicity). Ensemble averages at colliders have also been explored for $t\bar{t}$ production at the LHC, where spin correlations in dileptonic decays may be used both to detect entanglement and to perform quantum tomography for the $t\bar{t}$ pair [68]. More recently, polarisations reconstructed from scattering kinematics have been employed to test another fundamental feature of quantum theory, namely contextuality [69, 70]. Some of that work looks at polarisation for $W$-bosons created by (very short-timescale) top-quark decays, and therefore may have some bearing on the physics explored here.
@@ 4.1 Pure-state formula at fixed kinematics
In a highly idealised situation, namely a single kinematic configuration at fixed partonic centre-of-mass energy $\sqrt{s}$ and fixed angle between the tau and its parent quark, $\theta$, the spin state of the $\tau^{-}\tau^{+}$ pairs can be written in amplitude form
[…]
@@
Let it be clear that $\lvert$$\rangle$ refers to $\tau^{-}$ with helicity $\tfrac{1}{2}$ and $\tau^{+}$ with helicity $\tfrac{1}{2}$, whereas $\lvert-$$\rangle$ is the opposite. Thus, $\mathcal{M}_{+}$ corresponds to $\tau^{-}$ having helicity $\tfrac{1}{2}$, and $\mathcal{M}_{-}$ corresponds to $\tau^{-}$ having helicity $\tfrac{1}{2}$. No summation over angles or energies is included. In a perfect one-angle, one-energy scenario, such as an $e^{+}e^{-}$ collision at a fixed $\sqrt{s}$ and fixed $\theta$, this amplitude-level expression is essentially the mean helicity of the sub-process.
@@ 4.2 Ensemble-level cross-section formula
where $\sigma_{+}$ and $\sigma_{-}$ denote the cross sections for producing $\tau^{-}$ with helicity $\tfrac{1}{2}$ and $\tfrac{1}{2}$, respectively. In density operator language, each configuration corresponds to a pure spin state, whereas the final state is an overall mixture. Because hadron colliders typically sample many angles and energies, this cross-section ratio is the quantity we actually observe. [^12] Many authors simply write $\mathcal{P}_{\tau}$ for the ratio (63). Strictly speaking, it is still an average over fiducial phase space, so one could add angle brackets $\langle\cdots\rangle$. However, the shorthand without brackets is common in theory papers.
4.3 Extracting average tau helicity experimentally
@@
where $I$ is the incident flux and NN is the observed number of events within a given kinematic configuration. Counting events in different angular bins thus serves as a proxy for the underlying cross section.
In the absence of detector effects and assuming full access to the $\tau$ rest frame, [^13] In practice, $\tau$ rest frame kinematics are obtained probabilistically by reconstructing the missing neutrino momenta. Examples include the Missing Mass Calculator [72] (a likelihood scan over the kinematically allowed $\tau^{+}\tau^{-}$ phase space) and Point-Edge Transformer [73] generative models that infer the neutrino momenta from event-level observables. the two-body decay kinematics imply that the direction of the visible decay product is determined by the spin of the parent $\tau$, due to angular momentum conservation and the fixed spin states of the daughter particles. For $\tau^{-}\rightarrow\pi^{-}\nu_{\tau}$, the neutrino is always left-handed (with spin opposite to its momentum). Therefore
  • A $\tau^{-}$ with spin $\tfrac{1}{2}$ emits the $\pi^{-}$ along the spin direction ($\cos\theta>0$);
  • A $\tau^{-}$ with spin $\tfrac{1}{2}$ emits the $\pi^{-}$ opposite to the spin direction ($\cos\theta<0$).
Here, $\theta$ is the angle between the $\tau$ spin direction and the pion momentum. Even though we are considering the $\tau$ in its rest frame and helicity is not strictly defined, the pion direction remains directly correlated with the spin state of the $\tau$. In this way, the pion emission angle serves as an experimental proxy for helicity.
@@
where $N_{f}$ is the number of events in the forward region and $N_{b}$ is the number in the backward region. Because the helicity state is (ideally) in one-to-one correspondence with the emission direction of the pion, this ratio captures the asymmetry of the $\tau^{-}$ with respect to $\tfrac{1}{2}$ and $\tfrac{1}{2}$ helicity.
@@
4.4 Preliminary bound on deviations from the quantum mean
A microscopic model of the failure of the Born rule, in general, will give a correction to the mean polarisation. For tau leptons, this will imply an anomalous mean helicity
$\langle P_{\tau}\rangle=\langle P_{\tau}\rangle_{\mathrm{QT}}+\varepsilon,$ (67)
where $\langle P_{\tau}\rangle_{\mathrm{QT}}$ is the quantum-theoretical prediction and $\varepsilon$ is a small parameter. Existing ATLAS and CMS measurements of mean helicity for tau leptons, originating from decays $Z\!\to\tau^{+}\tau^{-}$, can then be interpreted as providing bounds on $\varepsilon$.
For events where the partonic centre-of-mass energy $\sqrt{s}$ is known to be close to the $Z$-boson mass, we have
$\langle P_{\tau}\rangle_{\mathrm{QT}}=-\frac{2v_{\tau}a_{\tau}}{v^{2}_{\tau}+a^{2}_{\tau}},$ (68)
where $v_{\tau}$ and $a_{\tau}$ are the effective neutral-current and axial-vector couplings for the $\tau$ lepton [65]. For the case $v_{\tau}\ll a_{\tau}$, we may write
$\langle P_{\tau}\rangle_{\mathrm{QT}}\approx-2\frac{v_{\tau}}{a_{\tau}}=-2\left(1-4\,\sin^{2}\theta_{W}^{\rm eff}\right),$ (69)
where $\theta_{W}^{\rm eff}$ is the effective weak mixing angle. Taking $\sin^{2}\theta_{W}^{\rm eff}=0.2315\pm 0.0002$ [74] we find
$\langle P_{\tau}\rangle_{\mathrm{QT}}=-0.1472\pm 0.0016.$ (70)
This can be compared with the CMS fit result
$\langle P_{\tau}\rangle_{\mathrm{meas}}=-0.144\pm 0.006\ \rm{(stat)}\pm 0.014\ \rm{(syst)}=-0.144\pm 0.015$ (71)
(where statistical and systematic uncertainties are reported separately and then combined). [^14] This result includes multiple $\tau$ decay modes. If one were to solely focus on the $\tau^{-}\to\pi^{-}\nu_{\tau}$ decay channel, one would get cleaner results at the expense of less statistics.
If we interpret the experiment as it stands as providing a constraint on $\varepsilon$, formally we find
$\varepsilon=0.003\pm\sigma_{\varepsilon},$ (72)
where the total uncertainty $\sigma_{\varepsilon}$ is at least 0.015, since
$\sigma_{\varepsilon}=\sqrt{\sigma_{\text{meas}}^{2}+\sigma_{\text{QT}}^{2}}\geq\sigma_{\text{meas}}=0.015.$ (73)
Here, $\sigma_{\text{meas}}$ is the uncertainty from the measurement of mean helicity and $\sigma_{\text{QT}}$ is the uncertainty in the quantum-theoretical prediction coming from the uncertainty in the effective angle $\theta_{W}^{\rm eff}$. If $\sigma_{\text{QT}}$ is neglected, we have a significance of only $0.2\sigma$. Including the uncertainty in $\theta_{W}^{\rm eff}$ would reduce the significance further. The experiment therefore clearly favours $\varepsilon=0$ and we can make no meaningful claim for a preferred $\varepsilon\neq 0$.
To detect a deviation of $\varepsilon=0.003$ at a $3\sigma$ level, we would need a total uncertainty of $\sigma_{\varepsilon}\approx 0.001$. The High-Luminosity LHC (HL-LHC) will have a luminosity increased by a factor of about $100$ compared to the cited CMS experiment, so the statistical uncertainty of $\langle P_{\tau}\rangle_{\mathrm{meas}}$ will be reduced by a factor of 10, from $\approx 0.006$ to about $0.0006$, which is already below our target. The limiting factor is therefore the systematic uncertainty, which in the cited CMS measurement is $\approx 0.014$.
More than half of the systematic uncertainty ($0.008$ out of $0.014$) arises from decay-mode migration, in particular the uncertainty in the migration matrix that maps true hadronically decaying $\tau$ modes to their reconstructed decay channels. To reach $\sigma_{\varepsilon}\approx 0.001$ at the HL-LHC, we would need a $\tau$-related systematic uncertainty $\sigma_{\mathrm{syst}}\approx 10^{-3}$, corresponding to an improvement by a factor of about 20 compared with the current analysis ($\sigma_{\mathrm{syst}}=0.014$). Achieving this reduction would require a significantly more accurate determination of $\tau$ decay modes, which is unlikely to be feasible without a focused experimental effort specifically optimised for $\tau$ polarisation.
@@ 5 Background assumptions and Born-rule deviations
In principle, there is clearly a tension when searching for deviations from the Born rule with an experiment whose analysis, in part, depends on the Born rule. However, in this paper we are concerned with testing the Born rule only as applied to simple two-state systems (spin or polarisation) whereas the experimental modelling and event selection applies the Born rule to multi-particle scattering amplitudes. On this point, there are two points to make. First, it could happen that the Born rule is violated only in the first case and not in the second case. Second, because multi-particle scattering observables are averaged over broad regions of phase space, it seems plausible that small deviations from the Born rule could be washed out, leaving significant deviations only for the simple two-state systems. To address this point in future work, in principle we would need to model the entire experimental analysis, which does not seem feasible without some simplifying assumptions. Here we propose adopting a preliminary working hypothesis: we assume that the Born rule holds for background modelling and event selection, while allowing for possible deviations only in the two-state polarisation observables.
@@ 6 Conclusion
We have discussed how certain experiments at high-energy colliders might be repurposed as tests of the Born rule on timescales as short as $\sim 10^{-25}\,\mathrm{s}$. In ideal conditions, measuring the polarisation of individual gamma-ray photons would enable us to probe polarisation probabilities as a function of angle, providing detailed constraints on Born-rule violations, as parameterised here in terms of expectation values $E(\mathbf{m})$ that are nonlinear in the Bloch measurement axis $\mathbf{m}$. However, for such high-energy photons, current technology appears to permit measurement of average polarisation only, limiting the search for Born-rule violations to anomalies in the effective polarisation vector $\mathbf{P}$, which need not coincide with the usual quantum vector $\mathbf{P}_{\mathrm{QT}}$ as defined by (5). We have illustrated this by measurements of mean helicity for tau leptons, enabling us to set a simple preliminary bound on Born-rule violations.
@@
Acknowledgements. We warmly thank Oliver K. Baker for his interest in and support of this work. MV is also grateful to Paul L. Tipton for his helpful advice while revising this paper. MV is supported by the DOE Office of High Energy Physics under Grant No. DE-SC0017660.